# Deep learning interpretability analysis for carbon star identification in DR3

### v1

@@ 3.1 GaiaNet
Convolutional Neural Network (CNN; LeCun etΒ al. 1998) is proposed by the concept of perceptual field in biology. By setting several convolutional kernels of different sizes, the model can effectively capture a wider range of global and local information, thereby enhancing the feature extraction of the input signal. The neural network model proposed in this work is tentatively named βGaiaNetβ.
@@ 3.1.1 The structure of GaiaNet
There is experimental evidence that extending the depth of a 1D convolutional model can effectively enhance the modelβs fitting and feature learning capabilities (Chen 2015). A parallel CNN structure consisting of multiple sets of convolutional kernels of different sizes can effectively capture key features of different sizes and achieve similar effects to ensemble learning. Both the serial deep structure and the wider parallel structure scheme can effectively capture key features. We have drawn on these advantages, which are well reflected in the improvement of our model. We first modified the size of the convolutional kernel in the input layer so that it can receive one-dimensional data, i.e., data of shape 1 Γ 343. We added regularization units consisting of BN (Batch normalization; Ioffe & Szegedy 2015) and Dropout. Both methods are effective in improving the generalization of the model. For the data in each batch, the BN layer first normalizes them. The formula for BN is as following:
@@
where denotes the original output of neuron and denotes the output of neuron after Dropout. During the training phase, each neuron is randomly selected and turned off with a probability of and retained with a probability of . The retained neuron needs to be scaled by dividing its output value by to keep the expectation of the output value constant. We add Dropout before the final fully connected layer of the model to prevent overfitting and to improve the generalization of the model.
Our model is made up of 16 parallel network structures, each with different-sized convolutional input layers to be able to better capture and extract key features in different fields of view. The same deep hidden layer network structure is then used, which consists of three convolutional layers, two pooling layers, a maximum pooling layer, and a average pooling layer. The deeper hidden layer structure helps to enhance the modelβs non-linear fitting capability, allowing for the extraction of key non-linear features. The overall architecture of the model is derived from βtextCNNβ, which has been improved by several iterations. We referenced the NIN network by introducing a 1Γ1 convolution, and a global average pooling operation (Lin etΒ al. 2013). The Inception Module network structure from GoogLeNet is borrowed, i.e., a structure consisting of multiple parallel and smaller dense convolutions with pooling approximating sparse connections (Szegedy etΒ al. 2015). The flow chart of the model is as Figure 6.
The 1Γ1 convolution (Lin etΒ al. 2013) can further deepen the network, effectively reduce the number of parameters in the overall convolution layer of the model, enhance the modelβs non-linear fitting ability, allow for flexible dimensioning up and down of the feature map, and enable cross-channel information interaction and integration. Global averaging pooling replaces the original practice of stitching the maximum pooled feature maps directly into the fully connected layer. These feature maps are passed through the global average pooling layer and then concatenated to the linear layer to give the corresponding categories. The purpose of this is to share information from the multi-channel feature maps so that they can all contribute to the final result, which we believe makes more sense and is one of the main innovations of this model. The formula is as below:
The global average pooling is handled very simply, where denotes the feature map value at the -th row position of the -th input tensor and denotes the height of the tensor. The global averaging pooling layer therefore calculates the average of all pixel points within a single tensor activation feature map. The overall structure of the model is shown in the Figure 7.
@@ 3.1.2 Important model parameters
The output layer of our model finally passes through a Sigmoid activation function (Lippmann 1987), which is an βSβ-type function that converts the real values of the output into a 0-1 probability distribution to represent the confidence level of a positive case. We usually set 0.5 as the threshold value, cases greater than the threshold being classified as positive, and cases less than or equal to the threshold being classified as negative. It is in line with the 0,1 label defined when we process secondary classification tasks. We use the BCELoss (Binary Cross Entropy Loss; Krogh & Hertz 1991) function to measure the difference between the model output and the true label, which can be interpreted as a maximum likelihood estimate, i.e. maximizing the probability of the observed data and finally converging the loss by an optimizer. It is a natural pairing with the Sigmoid activation function when dealing with dichotomous problems, which makes the model output easier to interpret and understand. The formula for calculating the loss of the output from the Sigmoid activation function using BCELoss is as below:
@@
Furthermore, the batch_size is the number of input cases in a batch used to update the model parameters for each iteration. A larger batch_size can speed up convergence by reducing the number of parameter updates. A smaller batch_size can improve the generalization ability of the model. By randomly selecting small batches of case for parameter updates introduces greater variability in each batch, facilitating the modelβs search for a global optimal solution without overfitting the training set. The selection of batch_size should be comprehensively considered based on factors such as the size of the dataset, the nature of the task, memory and computing power.
@@
Finally, there are several parameters that play a crucial role in the SGD (Stochastic Gradient Descent) optimization algorithm during the training phase of the model. The (learning rate) is a factor that controls the step size of the model weights update, which represents the size of the updated weights at each iteration. A large learning rate may lead to fast convergence of the model during training, but may also lead to the model skipping the optimal solution. A smaller learning rate may result in slower convergence of the model, but it is more likely to achieve a better optimal solution. In our practical tests, we opted for a smaller learning rate to avoid missing the global optimal solution. During the SGD training shown in Figure 8, the size of each step is fixed, but with the introduction of the momentum learning algorithm, the movement of each step depends not only on the magnitude of the current gradient but also on the accumulation of past velocities. Momentum uses historical gradient information to adjust the direction and speed of parameter updates (Polyak 1964), thus accelerating the convergence process of the SGD gradient descent and reducing oscillations during the gradient update process. It can be seen as an inertia introduced to leverage the previous gradient information in parameter updates. The core idea of momentum is to introduce an accumulated gradient history variable, which is similar to momentum in physics and records the direction and velocity of the previous gradientβs motion. During each iteration of the update, the momentum algorithm considers not only the current gradient but also the trends of previous gradients. This allows the parameters to be updated with a certain βmomentumβ in the direction of the gradient, resulting in faster crossing of flat areas and less invalid oscillation. The parameter update formula in the momentum algorithm is shown below:
where ΒΏ 0 represents the learning rate, is the momentum parameter, and is the gradient at (Sutskever etΒ al. 2013). We were surprised to find in our practical tests that it was excellent for improving model accuracy and further converging model loss.
@@ 3.2 SHAP
Specifically, for a deep learning model , where denotes the input features and denotes the class or real value of the output, is the number of features. The explanation model matches the original model when , where represents the model output with all simplified inputs toggled off (i.e. missing). In a word, SHAP defines the Shapley value for an input as , which represents the marginal contribution of feature to the model prediction.
@@
This method allows us to compare the Shapley values of different features to determine how much they affect the output of the model. It has been flexibly applied to various deep learning models, including convolutional neural networks (CNN), recurrent neural networks (RNN; Werbos 1988), and others. We will use SHAP in our model to interpret the key features that affect the model output. In addition, SHAP provides several visualization tools and methods, which will help us understand the model better.
@@ 3.3 Evaluation index
In addition to the accuracy criterion, we also introduced more refined evaluation metrics such as recall and precision to effectively measure the modelβs data mining performance during dataset validation. We define TP (true positive) as the number of cases whose categories are true and predicted categories are also positive, FN (false negative) as the number of cases whose categories are true but predicted categories are negative, FP (false positive) as the number of cases whose categories are false but predicted categories are positive, and TN (true negative) as the number of cases whose categories are false and predicted categories are also negative.
Accuracy refers to the proportion of the number of accurate cases classified by all categories to the total number of cases. We can calculate it as below:
@@
We also use confusion matrices to see the performance of the model more intuitively, and confusion matrices provide a good indication of the distribution of the various cases. For βs Carbon Star data mining work, the idea is to have as many TP cases as possible and as few FN cases as possible, i.e., as large a recall rate as possible, so we use this as our main validation set measure.
@@ 4 Results
4.1 Analysis and experiments of Dataset
Almost all C-N type carbon stars are in the AGB evolution stage, of which the carbon abundances in their atmosphere were produced in the stellar interiors, and were brought to the surface by the third dredge-up process. Almost all of them are intrinsic, thermally pulsing AGB stars, i.e., TP-AGB stars with large luminosities (Abia etΒ al. 2002). This kind of carbon stars typically exhibit stronger C and CN molecular bands than usual stars with temperatures below 3800βK (Gaia Collaboration etΒ al. 2023). Of course, there are also some kinds of carbon stars with relatively high surface temperatures, whose spectral types can be K or even G, and they are often extrinsic carbon stars. However, there is a large number of giant stars with the spectral type G, K, and M, which have similar colors, surface temperatures, and luminosities to the carbon stars, making it impossible to distinguish between the carbon stars and other giants just from the HR diagrams, but now possible through βs spectroscopy. Lebzelter etΒ al. (2023) and Sanders & Matsunaga (2023) have demonstrated its potential for separating C-rich and O-rich stars.
@@ 4.1 Analysis and experiments of Dataset
After the extinction correction, we calculate the absolute magnitudes and the intrinsic color index of for 4002 positive sample stars, 6,761 negative sample stars, and 64,761 stars of CSTAR_XPM sample, respectively. Most of the CSTAR_XPM stars have 1 Β‘ Β‘ 3 and Β‘ 2 mag, and most of the carbon stars have 2 Β‘ Β‘ 4 and Β‘ -1 mag. This suggests that the CSTAR_XPM_G sample may consist mostly of brighter and cooler C-N type carbon stars. From the corrected color-magnitude diagram, we also found that about half of the negative sample stars have a , while the of M giants should usually be larger than 1.6. We estimated that there are about half of the G and K giants in the CSTAR_XPM_nonG, which is corroborated by our later cross-matched results with the LAMOST DR10 spectroscopic data in Section 5.1. Hence, the CSTAR sample also includes a large mixture of other types of stars, rather than overwhelmingly M stars. In addition, a portion of the CSTAR_XPM_G has Β‘ 2 and relatively lower luminosities. From Figure 11 of Li etΒ al. (2024), it can be inferred that they may be extrinsic warm carbon stars with spectral types of G or K like Ba stars. It can be seen that there is a slightly clearer distinction between positive sample and negative sample in the matted color magnitudes, but some of the stars are still mixed together. However, elemental abundances are also reflected in the spectra. Specifically, the carbon stars are characterized by the strong absorption lines of carbon molecules in their spectra. This offers the possibility of a spectroscopic data-driven approach to distinguish the two directly from the stellar spectra. C2023 proposed screening carbon stars by two CN molecular band head strengths (R and R), which could be partially contaminating through our inspection, and would result in the loss of some potential carbon stars.
As mentioned in Section 2.2, we carefully selected a negative sample of 8556 spectra from CSTAR_XPM_nonG and a positive sample of 8176 spectra from CSTAR_XPM_G. We equalized the number of positive and negative samples by randomly adding negative spectra using the resampling techniques. So we composed a training dataset after fine visual inspection and machine ranking to try to make the better distinction in a deep learning way and expect meaningful interpretable analysis.
@@ 4.1.2 Comparison with other algorithms
In the past decades, many machine learning includes integrated learning methods (KNN: Cover & Hart 1967; SVM: Cortes & Vapnik 1995; Random Forest: Breiman 2001; XGBoost: Chen & Guestrin 2016 etc.) have been successfully applied to stellar spectral classification and physical parameters estimation. In contrast, convolutional neural networks are well known to have been used with great success in the field of image recognition, so we see the potential for the use of one-dimensional convolutional deep learning networks in spectral recognition.
@@
Recall is a very important evaluation index that effectively measures the rate of retaining positive sample. Our model retains all stars of positive sample and demonstrates approximately a 0.43% improvement over other algorithms, resulting in an overall accuracy increase of about 0.3%. The reason for the error rate of 0 in the validation set is likely to be the excellent overall architecture of our model, which is mainly reflected in the natural advantages of the convolutional model in feature extraction, as well as some necessary innovative improvements. On the other hand, we have conducted βexpertβ screening and purification of the training data to exclude erroneous sprctra and ensure the high purity and reliability of the training data, combined with some necessary data enhancement methods to ensure that the model can learn the real key information. Of course, the spectra of golden sample almost all have strong molecular bands, making them easy to distinguish from the spectra in CSTAR_XPM_nonG.
@@ 4.2 Model interpretation and key features
The SHAP value is used to quantify the effect of features on the model output. We consider each wavelength data point in a spectrum as a feature, and the corresponding flux to each wavelength as the feature value. The SHAP interpretable model can explain the output of our deep learning model by calculating a SHAP value for each feature of a given spectrum, which indicates the importance impact or contribution of the feature to the model output (the confidence of a carbon star as mentioned in subsection 3.1.2). A positive SHAP value indicates that the feature contributes positively to the modelβs prediction of the spectrum being a carbon star, while a negative SHAP value indicates that the feature contributes negatively to the modelβs prediction of the spectrum being a carbon star. The larger the absolute value of SHAP, the greater the impact of the feature on the model output. For a spectrum, the sum of the SHAP values of each feature can be used to approximate the output of the classification model. We can also calculate the distribution of the global most important features based on the statistics of SHAP values from a large sample.
@@
To get an overview of which features are most important and intuitively understand the influence of them on the model prediction results, we randomly selected 1024 carbon stars of positive sample and 1024 non-carbon stars of negative sample as background spectra of the deep explainer interpreter of SHAP, calculated the SHAP values of these spectra, and plotted the beeswarm in Figure 12. The top 30 most important features were shown in this figure, which were ranking by the sum of absolute SHAP value magnitudes of them over all 2048 spectra, and uses SHAP values to show the distribution of the impacts each feature has on the model output. For a particular feature, SHAP first normalize its values across all samples, the normalized feature values are then mapped onto a color gradient bar. The feature point will be assigned different colors (from blue to red), the color blue represents the feature point value is relatively small, and red means that feature point value is relatively large.
@@
Therefore, the beeswarm plot can be very intuitive reflect the influence of the feature on the output results under different values. As shown in the figure, we can see that the wavelength range 8900-9000βΓ Β is the location of the peak arising from the molecular band head. The larger the values of these feature points are, the more positively the model predicts carbon stars. Near 8100βΓ Β and 9300βΓ Β are the locations of the troughs arising from the molecular absorption band. The smaller these feature values are, the more negatively the model predicts carbon stars. It is evident that the most striking distinguishing feature is the presence of these two troughs.
@@
To better illustrate the differences in key feature contributions within a single spectrum, the background spectrum for calculating SHAP values can be chosen as the smooth pseudo-continuum after polynomial fitting using the approach of Zhang etΒ al. (2021). It serves as a baseline and reference for individually calculating the SHAP value of each feature in each spectrum. We believe that contrasting with such a baseline can effectively highlight the trough and peak features in each spectrum. Subsequently, we can generate an intuitive and clear feature thermal distribution map for the spectrum shown in Figure 13, which facilitates observing the distribution locations of the key features concerned by our model.
@@
However, if we want to get a more statistically significant result and reach a definitive conclusion, we need statistical information on the SHAP values of a large enough sample, which is more revealing than above. For this purpose, we calculate the average SHAP values of each feature for all positive and negative samples. As shown in Figure 14, we can see that the interpretation ranges of positive features for positive sample are consistent with the CN features. It is mainly distributed in the ranges 9140-9440βΓ , 7900-8360βΓ , 8820-9060βΓ , and 7600-7820βΓ , showing that relatively large SHAP values were assigned to both peak and trough ranges in the CN molecular bands. However, compared with the positive carbon star sample, the average SHAP values in these regions of the negative sample are much smaller. In addition, the negative sample showed a very strong negative contribution in the range 8420-8900βΓ , which is the most significant distinguishing feature of this sample. We note that the average SHAP values of the positive sample is also negative in the 8420-8620βΓ Β region. We suspect that this is related to the absorption feature of CaβII (8484-8662βΓ ), which is considered by the model to be a negative feature. We can see that the location of the modelβs main concern is the strong CN molecular band feature at the red end of the spectrum, which is consistent with the actual interpretation in Section 20. Next, we will showcase these key features.
@@
The reason our model only identifies the prominent CN features in the red end of the positive sample spectra, while almost failing to detect the C features in the blue end, can be attributed to two factors. Firstly, the flux values in the blue end are too small relative to the entire spectrum, making the C features much more insignificant compared to the prominent CN features. Secondly, C2023 filters golden sample carbon stars in the cold (most of them - ΒΏ 2; Gaia Collaboration etΒ al. 2023) CSTAR sample based on the prominent CN features rather than the much weaker C features. This results in more than half of the targets in the golden sample of carbon stars having almost no C features. Therefore, the C features in Gaiaβs XP spectra are not distinguishing for the identification of cold carbon stars.
@@
We combined the model training and validation accuracy, as well as the changes in FP and FN in Section 4.1.1, and ultimately determined the training epoch number of the model as 16. We applied the model trained based on the dataset provided in Section 3 to the 74,740 stars of CSTAR_XPM_nonG, and 451 new carbon star candidates were obtained. The spectra of these stars exhibit relatively prominent CN molecular bands compared to typical giant stars but are weaker than most golden sample stars. Their spectra show reduced flux at the blue end and hardly any significant C2 features are visible. In Figure below 15, we show some spectra of these sampl stars.

### v2

@@ 3.1 GaiaNet
Convolutional Neural Network (CNN; LeCun etΒ al. 1998) is proposed by the concept of perceptual field in biology. By setting several convolutional kernels of different sizes, the model can effectively capture a wider range of global and local information, thereby enhancing the feature extraction of the input signal. The neural network model proposed in this work is named βGaiaNetβ.
@@ 3.1.1 The structure of GaiaNet
There is experimental evidence that extending the depth of a 1D convolutional model can effectively enhance the modelβs fitting and feature learning capabilities (Chen 2015). A parallel CNN structure consisting of multiple sets of convolutional kernels of different sizes can effectively capture key features of different sizes and achieve similar effects to ensemble learning. We have drawn on these advantages, which are well reflected in the improvement of our model. We first modified the size of the convolutional kernel in the input layer so that it can receive one-dimensional data, i.e., data of shape 1 Γ 343. We added regularization units consisting of BN (Batch normalization; Ioffe & Szegedy 2015) and Dropout. Both methods are effective in improving the generalization of the model. For the data in each batch, the BN layer first normalizes them. The formula for BN is as following:
@@
where denotes the original output of neuron and denotes the output of neuron after Dropout. During the training phase, each neuron is randomly selected and turned off with a probability of and retained with a probability of . The retained neuron needs to be scaled by dividing its output value by to keep the expectation of the output value constant. We add Dropout before the final fully connected layer of the model to prevent overfitting and to improve the modelβs generalization.
Our model is made up of 16 parallel network structures, each with different-sized convolutional input layers to be able to better capture and extract key features in different fields of view. The same deep hidden layer network structure is then used, which consists of three convolutional layers, two pooling layers, a maximum pooling layer, and a average pooling layer. The deeper hidden layer structure helps to enhance the modelβs non-linear fitting capability, allowing for the extraction of key non-linear features. The overall architecture of the model is derived from βtextCNNβ, which has been improved by several iterations. We referenced the NIN network by introducing a 1Γ1 convolution, and a global average pooling operation (Lin etΒ al. 2013). The Inception Module network structure from GoogLeNet is borrowed, i.e., a structure consisting of multiple parallel and smaller dense convolutions with pooling approximating sparse connections (Szegedy etΒ al. 2015). The flow chart of the model is presented in Figure 6.
The 1Γ1 convolution (Lin etΒ al. 2013) can further deepen the network, effectively reduce the number of parameters in the overall convolution layer of the model, enhance the modelβs non-linear fitting ability, allow for flexible dimensioning up and down of the feature map, and enable cross-channel information interaction and integration. Global averaging pooling replaces the original practice of stitching the maximum pooled feature maps directly into the fully connected layer. These feature maps are passed through the global average pooling layer and then concatenated to the linear layer to give the corresponding categories. The purpose of this is to share information from the multi-channel feature maps so that they can all contribute to the final result, which is one of the main innovations of this model. The formula is as below:
The global average pooling is handled with parameters denotes the feature map value at the -th row position of the -th input tensor, and denotes the height of the tensor. The global averaging pooling layer calculates the average of all pixel points within a single tensor activation feature map. The overall structure of the model is shown in the Figure 7.
@@ 3.1.2 Important model parameters
The output layer of our model finally passes through a Sigmoid activation function (Lippmann 1987), which is an βSβ-type function that converts the real values of the output into a 0-1 probability distribution to represent the confidence level of a positive case. We usually set 0.5 as the threshold value, case greater than the threshold being classified as positive, and case less than or equal to the threshold being classified as negative. It is in line with the 0,1 label defined when we process secondary classification tasks. We use the BCELoss (Binary Cross Entropy Loss; Krogh & Hertz 1991) function to measure the difference between the model output and the true label, which can be interpreted as a maximum likelihood estimate, i.e., maximizing the probability of the observed data and finally converging the loss by an optimizer. It is a natural pairing with the Sigmoid activation function when dealing with dichotomous problems, which makes the model output easier to interpret and understand. The formula for calculating the loss of the output from the Sigmoid activation function using BCELoss is as below:
@@
Furthermore, the batch_size refers to the number of input cases in a batch used to update the model parameters during each iteration. A larger batch_size can speed up convergence by reducing the frequency of parameter updates. Conversely, a smaller batch_size often improves the modelβs generalization ability by introducing more variability into each batch, thereby preventing overfitting to the training set. However, the choice of batch_size should be comprehensively considered based on factors such as dataset size, task complexity, memory limitations, and computational resources.
@@
There are several parameters that play a crucial role in the SGD (Stochastic Gradient Descent) optimization algorithm during the training phase of the model. The (learning rate) is a factor that controls the step size of the model weights update, which represents the size of the updated weights at each iteration. A large learning rate may lead to fast convergence of the model during training, but may also lead to the model skipping the optimal solution. A smaller learning rate may result in slower convergence of the model, but it is more likely to achieve a better optimal solution. In our practical tests, we opted for a smaller learning rate to avoid missing the global optimal solution. During the SGD training shown in Figure 8, the size of each step is fixed, but with the introduction of the momentum learning algorithm, the movement of each step depends not only on the magnitude of the current gradient but also on the accumulation of past velocities. βMomentumβ uses historical gradient information to adjust the direction and speed of parameter updates (Polyak 1964), thus accelerating the convergence process of the SGD gradient descent and reducing oscillations during the gradient update process. It can be seen as an inertia introduced to leverage the previous gradient information in parameter updates. The core idea of momentum is to introduce an accumulated gradient history variable, which is similar to momentum in physics and records the direction and velocity of the previous gradientβs motion. During each iteration of the update, the momentum algorithm considers not only the current gradient but also the trends of previous gradients. This allows the parameters to be updated with a certain βmomentumβ in the direction of the gradient, resulting in faster crossing of flat areas and less invalid oscillation. The parameter update formula in the momentum algorithm is shown below:
where ΒΏ 0 represents the learning rate, is the momentum parameter, and is the gradient at (Sutskever etΒ al. 2013). We found in our practical tests that it was excellent for improving model accuracy and further converging model loss.
@@ 3.2 SHAP
Specifically, for a deep learning model , where denotes the input features and denotes the class or real value of the output, is the number of features. The explanation model matches the original model when , where represents the model output with all simplified inputs toggled off (i.e. missing). SHAP defines the Shapley value for an input as , which represents the marginal contribution of feature to the model prediction.
@@
This method allows us to compare the Shapley values of different features to determine how much they affect the output of the model. It has been flexibly applied to various deep learning models, including convolutional neural networks (CNN), recurrent neural networks (RNN; Werbos 1988), and others. We will use the SHAP in our model to interpret the key features that affect the model output. In addition, SHAP provides several visualization tools and methods, which will help us understand the model better.
@@ 3.3 Evaluation index
In addition to the accuracy criterion, we also introduce other evaluation metrics such as recall and precision to effectively measure the modelβs data mining performance during dataset validation. We define TP (true positive) as the number of cases whose categories are true and predicted categories are also positive, FN (false negative) as the number of cases whose categories are true but predicted categories are negative, FP (false positive) as the number of cases whose categories are false but predicted categories are positive, and TN (true negative) as the number of cases whose categories are false and predicted categories are also negative.
Accuracy refers to the proportion of the number of accurate cases classified by all categories to the total number of cases. We calculate it as below:
@@
We also use confusion matrix to see the performance of the model more intuitively, as it effectively illustrate the distribution across different classes. For our carbon star data mining work, the idea is to have as many TP cases as possible and as few FN cases as possible, i.e., as large a recall rate as possible, so we use this as our main validation set diagnostic.
@@ 4 Results
4.1 Analysis and experiments on the dataset
Almost all N-type carbon stars are in the AGB evolution stage, where the carbon abundances in their atmosphere were produced in the stellar interiors, and were brought to the surface by the TDU process. Almost all of them are intrinsic, thermally pulsing AGB stars, i.e., TP-AGB stars with large luminosities (Abia etΒ al. 2002). This kind of carbon star typically exhibit stronger C and CN molecular bands than usual stars with temperatures below 3800βK (Gaia Collaboration etΒ al. 2023). Of course, there are also some kinds of carbon stars with relatively high surface temperatures, whose spectral types can be K or even G, and they are often extrinsic carbon stars. However, there is a large number of giant stars with the spectral type G, K, and M, which have similar colors, surface temperatures, and luminosities to the carbon stars, making it impossible to distinguish between the carbon stars and other giants just from the HR diagrams, but now possible through βs spectroscopy. Lebzelter etΒ al. (2023) and Sanders & Matsunaga (2023) have demonstrated its potential for separating C-rich and O-rich stars.
@@ 4.1 Analysis and experiments on the dataset
After the extinction correction, we calculate the absolute magnitudes and the intrinsic color index of for 4002 positive sample stars, 6761 negative sample stars, and 64,761 stars of CSTAR_XPM sample, respectively. Most of the CSTAR_XPM stars have 1 Β‘ Β‘ 3 and Β‘ 2 mag, and most of the carbon stars have 2 Β‘ Β‘ 4 and Β‘ -1 mag. This suggests that the CSTAR_XPM_G sample may consist mostly of brighter and cooler N-type carbon stars. From the corrected color-magnitude diagram, we also found that about half of the negative sample stars have a , while the of M giants should usually be larger than 1.6. We estimated that there are about half of the G and K giants in the CSTAR_XPM_nonG, which is corroborated by our later cross-matched results with the LAMOST DR10 spectroscopic data in Section 5.1. Hence, the CSTAR sample also includes a large mixture of other types of stars, rather than overwhelmingly M stars. In addition, a portion of the CSTAR_XPM_G has Β‘ 2 and relatively lower luminosities. If these sources are located in the crowded regions of the Galactic plane where the dust distribution is complex and non-uniform, the true interstellar medium does not follow a simple Gaussian process (Green etΒ al. 2019). In this case, their extinction must be treated with caution. If their color index estimates are correct, as suggested by Figure 11 of Li etΒ al. (2024), they are most likely extrinsic warm carbon stars with spectral types of G or K like Ba stars. It can be seen that there is a slightly clearer distinction between positive sample and negative sample in the color magnitudes, but some of the stars are still mixed together. However, elemental abundances are also reflected in the spectra (Abia etΒ al. 2002). Specifically, the carbon stars are characterized by the strong absorption lines of carbon molecules in their spectra. This offers the possibility of a spectroscopic data-driven approach to distinguish the two directly from the stellar spectra. C2023 proposed screening carbon stars by R and R, it could be partially contaminating and would result in the loss of some potential carbon stars through our inspection.
As mentioned in Section 2.2, we carefully selected a negative sample of 8556 spectra from CSTAR_XPM_nonG and a positive sample of 8176 spectra from CSTAR_XPM_G. We equalized the number of positive and negative samples by randomly adding negative spectra using the resampling techniques. So we composed a training dataset after fine visual inspection and machine ranking to try to ensure better distinction for deep learning way and meaningful interpretable analysis.
@@ 4.1.2 Comparison with other algorithms
In the past decades, many machine learning includes integrated learning methods (KNN: Cover & Hart 1967; SVM: Cortes & Vapnik 1995; Random Forest: Breiman 2001; XGBoost: Chen & Guestrin 2016 etc.) have been successfully applied to stellar spectral classification and physical parameter estimation. In contrast, convolutional neural networks are well known to have been used with great success in the field of image recognition, so we see the potential for the use of one-dimensional convolutional deep learning networks in spectral recognition.
@@
Recall is a very important evaluation index that effectively measures the rate of retaining positive sample. Our model retains all stars of positive sample and demonstrates approximately a 0.43% improvement over other algorithms, resulting in an overall accuracy increase of about 0.3%. The reason for the error rate of 0 in the validation set is likely to be the excellent overall architecture of our model, which is mainly reflected in the natural advantages of the convolutional model in feature extraction, as well as some necessary innovative improvements. On the other hand, we have conducted βexpertβ screening and purification of the training data to exclude erroneous spectra and ensure the high purity and reliability of the training data, combined with some necessary data enhancement methods to ensure that the model can learn the appropriate information. Though of course, The spectra of positive sample exhibit strong CN molecular bands, making them easily distinguishable from negative sample.
@@ 4.2 Model interpretation and key features
The SHAP value is used to quantify the effect of features on the model output. We consider each wavelength in a spectrum as a feature, and the corresponding flux to each wavelength as the feature value. The SHAP interpretable model can explain our deep learning model by calculating a SHAP value for each feature of a given spectrum, which indicates the importance impact or contribution of the feature to the model output (i.e. the confidence of a carbon star as mentioned in subsection 3.1.2). A positive SHAP value indicates that the feature contributes positively to the modelβs prediction of the spectrum being a carbon star, while a negative SHAP value indicates that the feature contributes negatively to the modelβs prediction of the spectrum being a carbon star. The larger the absolute value of SHAP, the greater the impact of the feature on the model output. For a spectrum, the sum of the SHAP values of each feature can be used to approximate the output of the classification model. We can also calculate the distribution of the most important features based on the statistics of SHAP values from a large sample.
@@
To get an overview of which features are most important and intuitively understand the influence of them on the model prediction results, we randomly selected 1024 carbon stars of positive sample and 1024 non-carbon stars of negative sample as background spectra of the deep explainer interpreter of SHAP, calculated the SHAP values of these spectra, and plotted in Figure 12. The top 30 most important features were shown in this figure, which were ranking by the sum of absolute SHAP value magnitudes of them over all 2048 spectra, and uses SHAP values to show the distribution of the impacts each feature has on the model output. For a particular feature, SHAP first normalizes its values across all spectra, the normalized feature values are then mapped onto a color gradient bar. The feature point will be assigned different colors (from blue to red), the color blue represents a feature point value that is relatively small, and red means that a feature point value is relatively large.
@@
Therefore, the beeswarm plot can be an intuitive reflection of the influence of the feature on the output results for different values. As shown in the figure, we can see that the wavelength range 8900-9000βΓ Β is the location of the peak arising from the molecular band head. The larger the values of these feature points are, the more positively the model predicts carbon stars. Near 8100βΓ Β and 9300βΓ Β are the locations of the troughs arising from CN molecular absorption bands. The smaller these feature values are, the more negatively the model predicts carbon stars. It is evident that the most striking distinguishing feature is the presence of these two troughs.
@@
To better illustrate the differences in key feature contributions within a single spectrum, the background spectrum for calculating SHAP values can be chosen as the smooth pseudo-continuum after polynomial fitting using the approach of Zhang etΒ al. (2021). It serves as a baseline and reference for individually calculating the SHAP value of each feature in each spectrum. We believe that contrasting with such a baseline can effectively highlight the trough and peak features in each spectrum. Subsequently, we can generate an intuitive and clear feature distribution heatmap for given spectrum shown in Figure 13, which facilitates the easy location of the key spectral features concerned by our model.
@@
However, if we want to get a more statistically signigicant and reach a quantified conclusion, we need statistical information on the SHAP values of a large enough sample. For this purpose, we calculate the average SHAP values of each feature for all positive and negative samples. As shown in Figure 14, we can see that the interpretation ranges of positive features for the positive sample are consistent with the CN features. It is mainly distributed in the ranges 9140-9440βΓ , 7900-8360βΓ , 8820-9060βΓ , and 7600-7820βΓ , showing that relatively large SHAP values were assigned to both peak and trough ranges in the CN molecular bands. However, compared with the positive carbon star sample, the average SHAP values in these regions of the negative sample are much smaller. In addition, the negative sample showed a very strong negative contribution in the range 8420-8900βΓ , which is the most significant distinguishing feature of this sample. We note that the average SHAP values of the positive sample is also negative in the 8420-8620βΓ Β region. We suspect that this is related to the absorption feature of CaβII (8484-8662βΓ ), which is considered by the model to be a negative feature. We can see that the location of the modelβs main concern is the strong CN molecular band feature at the red end of the spectrum, which is consistent with the actual interpretation in Section 20. Next, we will showcase these key features.
@@
The reason our model only identifies the prominent CN features in the red end of the positive sample spectra, while almost failing to detect the C features in the blue end, can be attributed to two factors. Firstly, the flux values in the blue end are too small relative to the entire spectrum, making the C features much less significant compared to the prominent CN features. Secondly, C2023 filtered golden sample carbon stars in relatively cold (most of them have - ΒΏ 2; Gaia Collaboration etΒ al. 2023) CSTAR sample based on the prominent CN features rather than the much weaker C features. This results in more than half of the targets in the golden sample of carbon stars having almost no C features.
@@
We combined the model training and validation accuracy, as well as the changes in FP and FN in Section 4.1.1, and ultimately determined the training epoch number of the model as 16. We applied the model trained based on the dataset provided in Section 3 to the 74,740 stars of CSTAR_XPM_nonG, and 451 new carbon star candidates were obtained. The spectra of these stars exhibit relatively prominent CN molecular bands compared to typical giant stars but are weaker than most golden sample stars. Their spectra show reduced flux at the blue end and hardly any significant C2 features are visible. In Figure 15 we show some spectra of these stars.